# Animate Your Thoughts: Decoupled Reconstruction of Dynamic Natural Vision from Slow Brain Activity

### v1

@@ Abstract
Reconstructing human dynamic vision from brain activity is a challenging task with great scientific significance. The difficulty stems from two primary issues: (1) vision-processing mechanisms in the brain are highly intricate and not fully revealed, making it challenging to directly learn a mapping between fMRI and video; (2) the temporal resolution of fMRI is significantly lower than that of natural videos. To overcome these issues, this paper propose a two-stage model named Mind-Animator, which achieves state-of-the-art performance on three public datasets. Specifically, during the fMRI-to-feature stage, we decouple semantic, structural, and motion features from fMRI through fMRI-vision-language tri-modal contrastive learning and sparse causal attention. In the feature-to-video stage, these features are merged to videos by an inflated Stable Diffusion. We substantiate that the reconstructed video dynamics are indeed derived from fMRI, rather than hallucinations of the generative model, through permutation tests. Additionally, the visualization of voxel-wise and ROI-wise importance maps confirms the neurobiological interpretability of our model.
@@ 1 Introduction
Researchers in computational neuroscience and the field of artificial intelligence have long sought to decipher and simulate the brain’s visual information processing mechanisms to advance the development of brain-inspired models [1, 2, 3].
In recent years, functional magnetic resonance imaging (fMRI) has emerged as a reliable tool for measuring brain activity due to its high spatial resolution as a non-invasive brain signal recording technique [4]. fMRI-based neural decoders, which map brain signals to visual stimuli, facilitate a deeper understanding of the human visual perception system. Neural decoding can be categorized into classification, identification, and reconstruction, with this study focusing on the most challenging aspect: reconstruction. Prior research has made significant strides in the classification [3, 5, 6, 7, 8] and identification [4, 9, 10, 11] of static stimulus images, and remarkably, some researchers have advanced to the point where they can reconstruct [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] images from brain signals that closely resemble the original stimulus images.
In reality, the majority of visual stimuli we encounter in daily life are continuous and dynamic. As depicted in Figure 1, when a subject views dynamic scenes, the primary visual cortex firstly processes low-level structural information like location, shape, size, and color [23], leading to the preliminary recognition of a black silhouette at the edge of a yellow background. Subsequently, motion information of the object is perceived [24], noting that the silhouette is moving from right to left. Lastly, in the higher visual cortex, the interpretation of category and description gives rise to high-level semantic understanding [25], comprehending the scene as a soldier walking from right to left in a desert.
@@
Due to the inherent nature of fMRI, which relies on the slow blood oxygenation level dependent (BOLD) [26, 27]signal, the sampling frequency is restricted to around 0.5Hz. This frequency is notably lower than the typical 30Hz frame rate of most videos. As a result, a significant discrepancy exists between the temporal resolution of fMRI and the nature video. In fact, each fMRI signal integrates information from approximately 60 video frames. This disparity makes the task of reconstructing video from fMRI signals an exceedingly complex challenge.
To address this challenge, Nishimoto [28] transforms the video reconstruction task into a identification problem, employing the Motion-Energy model [29] and Bayesian inference to reconstruct videos from a predefined video library. Subsequently, Han [30] and Wen [31] et al. map brain responses to the feature spaces of deep neural network (DNN) to reconstruct down-sampled (with the frame rate reduced to 1Hz) video stimuli. Wang [32] et al. develope an f-CVGAN that learns temporal and spatial information in fMRI through separate discriminators [33]. To mitigate the scarcity of video-fMRI data, Kupershmidt [34] et al. utilize self-supervised learning [35] to incorporate a large amount of unpaired video data. These efforts have validated the feasibility of video reconstruction from fMRI, albeit with a lack of explicit semantic information in the results. Chen [36] et al. utilize contrastive learning to map fMRI to the Contrastive Language-Image Pre-Training (CLIP) [37] representation space and fine-tuned inflated Stable Diffusion [38, 39] on a video-text dataset as a video generation model, successfully reconstructing coherent videos with clear semantic information for the first time. However, this work does not consider structure information such as color and position, and it is uncertain whether the motion information in the reconstructed videos originated from the fMRI or the video generation model.
Method Semantic Structure Motion Frame rate Resolution
Nishimoto [28] (Current Biology 2011) × × —— ——
Wen [31] (Cerebral Cortex 2017)
× × —— 64x64
Han [30] (NeuroImage 2019)
× × —— 128x128
Kupershmidt [34] × (✓) 4Hz 112x112
Wang [32] (Cerebral Cortex 2022)
× 4Hz 64x64
Chen [36] (NeurIPS 2023 Oral)
× (✓) 3Hz 256x256
Ours 4Hz 512x512
Table 1: Comparison of modal information used in Mind-Animator and related works. Parentheses indicate the utilization of external video data in the decoding of this feature.
@@
    As shown in Table 1, they fail to simultaneously capture semantic, structure, and motion information within the reconstructed videos. Moreover, the resolution of the video is low.
@@
    The reliance on external video datasets and video generation models introduces uncertainty regarding the true source of motion information, leading to the possibility that the reconstructed videos may represent a "hallucination" of the video generation model rather than an accurate dynamic decoding from the fMRI data.
@@
To address the aforementioned issues, we introduce Mind-Animator, a video reconstruction model that decouples semantic, structure, and motion information from fMRI, as illustrated in Figure 2. Specifically, we map fMRI to the CLIP representation space and the Vector Quantized-Variational Autoencoder (VQ-VAE) [40] latent space to capture semantic and structure information. We design a Transformer-based [41] motion decoder to extract motion information frame by frame from fMRI through a next-frame-prediction task. Finally, the decoded semantic, structure, and motion information is fed into an inflated Stable Diffusion [38, 39] without any fine-tuning to generate each frame of the video, ensuring that all information is derived solely from the fMRI data.
The contributions are summarized as follows:
  • (1)
@@
    We propose Mind-Animator, which for the first time successfully decouples semantic, structure, and motion information from fMRI to enable video reconstruction. To extract the motion and spatial information from fMRI, we propose temporal and spatial attention modules respectively, which decode subtle but significant motion information.
  • (2)
    We validate through a permutation test that the motion information in our reconstructed videos indeed originates from the fMRI, rather than being a "hallucination" generated by the video generation model.
  • (3)
@@
    We introduce seven evaluation metrics that comprehensively assess the reconstruction results of our model and all previous models across three dimensions—semantic, structure, and spatiotemporal consistency—on three publicly available high-quality video-fMRI datasets. Our model achieves state-of-the-art (SOTA) performance in five of these metrics and secures second place in the remaining two, with a notable 76% improvement in Structural Similarity Index (SSIM) over the previous SOTA. This establishes our work as the first comprehensive and unbiased benchmark for subsequent researchers. The code and data have been anonymously released at: https://github.com/Zuskd/Mind-Animator.
@@ 2 Methodology
Figure 2: The overall architecture of Mind-Animator, a two-stage video reconstruction model based on fMRI. As illustrated in subfigures (a), (b), and (c), three decoders are trained during the fMRI-to-feature stage to disentangle semantic, structural, and motion information from fMRI, respectively. Subfigure (d) demonstrates that, in the feature-to-video stage, the decoded information is input into an inflated Text-to-Image (T2I) model for video reconstruction.
2.1 Overview
Figure 2 presents the overall architecture of the proposed Mind-Animator, a video reconstruction model based on fMRI. The model consists of two stages: fMRI-to-feature and feature-to-video.
In the fMRI-to-feature stage, as depicted in Figure 1, we begin by emulating the human visual system’s approach to interpreting dynamic visual stimuli. This process involves the decomposition of video stimuli into high-level semantic feature, low-level structural feature, and motion feature. Then three separate decoders are trained to decode these features from fMRI: (a) for decoding semantic feature, we employ a contrastive learning loss to map fMRI into the visual-linguistic embedding space of CLIP[37], (b) we utilize the frame token extracted by VQ-VAE[40] as the video’s structural feature[42], followed by a simple Multi-Layer Perceptron (MLP) to fit it, and (c) we design a Transformer-based Consistency Motion Generator for decoding motion information. After training with a next-frame-prediction task, this module sequentially generates each subsequent frame token based on the first frame token decoded in section (b).
In the feature-to-video stage, depicted in Figure 2 (d), the decoded features are input into an inflated Text-to-Image (T2I) model, facilitating the reconstruction of the stimulus video without the interference of external training videos.
@@ 2.2 Problem Statement
We aim to decode videos from brain activity recorded with fMRI when healthy participants watch a sequence of natural videos. Let $X$ and $Y$ denote the voxel space and pixel space, respectively. Let $X_{i}\in R^{1\times n}$ be the fMRI signal when a video $V_{i,j}\in R^{1\times 3\times 512\times 512}$ is presented to the participant, where $n$ is the number of fMRI voxels, $j$ is the frame ID of video $i$ and $i\in[1,N]$, $j\in[1,8]$, with $N$ the total number of videos. Let $Z(k)$ denote the feature space, $k\in\{semantic,structure,motion\}$. The goal of fMRI-to-feature stage is to train decoders $D(k):X\rightarrow Z(k)$, and the goal of feature-to-video stage is to construct a video generation model $G:Z(semantic)\times Z(structure)\times Z(motion)\rightarrow Y$, without introducing motion information from external video data.
@@
Semantic Decoder
Due to the low signal-to-noise ratio of the fMRI signal $X_{i}$ and the substantial dimension discrepancy with the text condition $c_{i}\in R^{1\times 20\times 768}$, directly learning a mapping between them is prone to overfitting. Considering the robust semantic information embedded in the latent space of CLIP[43], and given that CLIP has been shown to outperform various single-modal DNNs in explaining cortical activity[44, 45], we employ bidirectional InfoNCE loss to align the fMRI with the latent space of CLIP (Vit-B/32)$\in R^{512}$, followed by a two-layer MLP to map it to text condition $c_{i}$,
$L_{BiInfoNCE}=-\frac{1}{B}\sum_{i=1}^{B}\Big{(}log\frac{exp(s(\hat{z}_{i},z_{i})/\tau)}{\sum_{j=1}^{B}exp(s(\hat{z}_{i},z_{j})/\tau)}+log\frac{exp(s(\hat{z}_{i},z_{i})/\tau)}{\sum_{k=1}^{B}exp(s(\hat{z}_{i},z_{k})/\tau)}\Big{)}.$ (1)
where $s$ is the cosine similarity, $z_{i}$ and $\hat{z}_{i}$ are the latent representation from two modalities, $B$ is the batch size, and $\tau$ is a learned temperature parameter. Then, given $f\in R^{B\times 512}$, $v\in R^{B\times 512}$, and $t\in R^{B\times 512}$ as the respective representations of fMRI, video, and text embeddings, the fMRI-vision-language trimodal loss is:
$L_{Semantic}=\alpha\cdot L_{BiInfoNCE}(f,t)+(1-\alpha)\cdot L_{BiInfoNCE}(f,v).$ (2)
Subsequently, to map the fMRI embedding $f_{i}$ to the text condition $c_{i}$ for the purpose of conditioning generative image models, a projection loss is utilized,
$L_{Projection}=\frac{1}{B}\sum_{i=1}^{B}\|MLP(f_{i})-c_{i}\|^{2}_{2}.$ (3)
@@
$L_{Combined}=\lambda_{1}\cdot L_{Semantic}+\lambda_{2}\cdot L_{Projection}.$ (4)
Structure Decoder
For a short video clip, it can be assumed that the low-level information (size, shape, and color) contained in each frame remains largely consistent with that of the first frame. Consequently, we utilize the token extracted from the first frame by VQ-VAE as structural information and train the structural decoder using the standard mean squared error (MSE) loss function. Let $\Phi$ denote the encoder of VQVAE, the structure loss is defined as:
$L_{Structure}=\frac{1}{B}\sum_{i=1}^{B}\|D_{Structure}(f_{i})-\Phi(V_{i,1})\|^{2}_{2}.$ (5)
Consistency Motion Generator
Inspired by natural language processing, we treat each video frame token as a word embedding, and develop an L-layer Transformer-based Consistency Motion Generator. For a more detailed introduction, please refer to Appendix B.1.
Figure 3: The detailed architectural diagram of the Consistency Motion Generator.
In the Temporal Module, visible video frame tokens $\Phi(V_{i})\in R^{m\times d_{token}}$ and positional encoding $E_{pos}\in R^{m\times d_{token}}$ are jointly input into a Sparse Causal Self-Attention (SCSA) layer to learn inter-frame temporal information. This attention layer incorporates a specially designed Sparse Causal Mask to ensure sparsity between frames and accelerate training. As illustrated in Figure 3, the mask is divided into fixed and random components. The fixed mask ensures that each frame cannot access information from subsequent frames, while the random mask maintains sparsity among visible frames, preventing the model from taking shortcuts[52]. During inference, we eliminate the random mask. As shown in Eq. 2.3, the model also applies residual connections and layer normalization (LN) to the variable $z_{l}$,
$z_{0}=$ $[\Phi(V_{i,1}),\Phi(V_{i,2}),\ldots,\Phi(V_{i,m})]+E_{pos},$
$z_{l}=$ $LN(SCSA(z_{l-1}))+z_{l-1}.\quad l=1,2,\ldots,L$ (6)
As shown in Eq. 7, in the Spatial Module, the embedding of the visible frames $z_{l}$ serves as the Query, while the fMRI signal $f$, after passing through an embedding layer, serves as the Key and Value in the cross-attention block. Following residual connections and layer normalization, $z_{l}$ is input into the Feed Forward Network (FFN) to predict the subsequent unseen frame tokens $\hat{\Phi(V_{i,j})},j\in[m+1,8]$:
$z_{l}=$ $CrossAttention(Q,K,V),\quad l=1,2,\ldots,L$ (7)
$Q=$ $W_{Q}^{l}\cdot z_{l},\quad K=W_{K}^{l}\cdot Emb(f),\quad V=W_{V}^{l}\cdot Emb(f),$
$z_{l}=$ $FFN(LN(z_{l})+z_{l-1}).\quad l=1,2,\ldots,L$ (8)
@@
$L_{Consistency}=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=m+1}^{8}\|\hat{\Phi(V_{i,j})}-\Phi(V_{i,j})\|^{2}_{2}.$ (9)

### v2

# Animate Your Thoughts: Reconstruction of Dynamic Natural Vision from Human Brain Activity

@@ Abstract
Reconstructing human dynamic vision from brain activity is a challenging task with great scientific significance. Although prior video reconstruction methods have made substantial progress, they still suffer from several limitations, including: (1) difficulty in simultaneously reconciling semantic (e.g. categorical descriptions), structure (e.g. size and color), and consistent motion information (e.g. order of frames); (2) low temporal resolution of fMRI, which poses a challenge in decoding multiple frames of video dynamics from a single fMRI frame; (3) reliance on video generation models, which introduces ambiguity regarding whether the dynamics observed in the reconstructed videos are genuinely derived from fMRI data or are hallucinations from generative model. To overcome these limitations, we propose a two-stage model named Mind-Animator. During the fMRI-to-feature stage, we decouple semantic, structure, and motion features from fMRI. Specifically, we employ fMRI-vision-language tri-modal contrastive learning to decode semantic feature from fMRI and design a sparse causal attention mechanism for decoding multi-frame video motion features through a next-frame-prediction task. In the feature-to-video stage, these features are integrated into videos using an inflated Stable Diffusion, effectively eliminating external video data interference. Extensive experiments on multiple video-fMRI datasets demonstrate that our model achieves state-of-the-art performance. Comprehensive visualization analyses further elucidate the interpretability of our model from a neurobiological perspective. Project page: https://mind-animator-design.github.io/.
@@ 1 Introduction
Advances in sensory neuroscience offer new perspectives on brain function and could enhance artificial intelligence development (Palazzo et al. (2021); Yargholi & Hossein-Zadeh (2016)). One of the critical aspects to the research is neural decoding, which links visual stimuli to corresponding functional magnetic resonance imaging (fMRI) brain recordings. Neural decoding methods include classification, identification, and reconstruction, with this study focusing on the most challenging aspect: reconstruction.
Prior methods have made significant strides in the classification (Yargholi & Hossein-Zadeh (2016); Horikawa & Kamitani (2017); Fujiwara et al. (2013)) and identification (Kay et al. (2008); Wildgruber et al. (2005)) of static stimulus images. Remarkably, some researchers have advanced to the point where they can reconstruct (Naselaris et al. (2009); Van Gerven et al. (2010); Chen et al. (2023); Takagi & Nishimoto (2023); Ozcelik et al. (2022); Beliy et al. (2019)) images from brain signals that closely resemble the original stimulus images. In reality, the majority of visual stimuli we encounter in daily life are continuous and dynamic, hence there is a growing interest in reconstructing video from brain signals. Building on previous work that decoupled semantic and structural information from fMRI to reconstruct images (Scotti et al. (2024); Lu et al. (2023); Fang et al. (2020)), we argue that when the visual stimulus shifts from static images to dynamic videos, as shown in Figure 1, it is crucial to account for three dimensions: semantic, structural, and motion, considering the brain’s processing of dynamic visual information.
Due to the inherent nature of fMRI, which relies on the slow blood oxygenation level dependent (BOLD) signal (Logothetis (2002); Kim & Ogawa (2012)), neural activity is integrated over a period exceeding 10 seconds ( 300 video frames). This integration delay poses a fundamental challenge in capturing rapid motion dynamics. Consequently, the task of reconstructing videos from fMRI signals becomes exceedingly challenging.
@@
To address this challenge, Nishimoto et al. (2011) transforms the video reconstruction task into an identification task, employing the Motion-Energy model (Adelson & Bergen (1985)) and Bayesian inference to retrieve videos from a predefined video library. Subsequently, Han et al. (2019), Wen et al. (2018) and Wang et al. (2022) map brain responses to the feature spaces of deep neural network (DNN) to reconstruct video stimuli. To mitigate the scarcity of video-fMRI data, Kupershmidt et al. (2022) utilizes self-supervised learning (Kingma & Welling (2014)) to incorporate a large amount of unpaired video data. While these efforts have confirmed the feasibility of video reconstruction from fMRI, the results are notably deficient in explicit semantic information. Chen et al. (2024) utilizes contrastive learning to map fMRI to the Contrastive Language-Image Pre-training (CLIP) (Radford et al. (2021)) representation space and co-training with a video generation model, successfully reconstructing coherent videos with clear semantic information for the first time. However, this work does not consider structure information such as color and position, and it is uncertain whether the motion information in the reconstructed videos originate from the fMRI or the external video data used in training the video generation model.
@@
    They fail to simultaneously capture semantic, structure, and motion information within the reconstructed videos.
@@
    The reliance on external video datasets and video generation models introduces ambiguity regarding whether the dynamics observed in the reconstructed videos are genuinely derived from fMRI data or are hallucinations from video generative model.
@@
To address the issues, we introduce Mind-Animator, a video reconstruction model that decouples semantic, structure, and motion information from fMRI, as illustrated in Figure 3. Specifically, we map fMRI to the CLIP representation space and the Vector Quantized-Variational Autoencoder (VQ-VAE) (Van Den Oord et al. (2017)) latent space to capture semantic and structure information. We design a Transformer-based (Vaswani et al. (2017)) motion decoder to extract motion information frame by frame from fMRI through a next-frame-prediction task. Finally, the decoded semantic, structure, and motion information is fed into an inflated Stable Diffusion (Rombach et al. (2022); Wu et al. (2023)) without any fine-tuning with video data to generate each frame of the video.
Our contributions are summarized as follows:
@@
(1) Method: We propose Mind-Animator, which enables video reconstruction by decoupling semantic, structural, and motion information from fMRI data for the first time.
To address the mismatch in timescales between fMRI and video data, we propose a Consistency Motion Generator with Sparse Causal Attention. This model decodes subtle yet significant motion patterns through a next-frame token prediction task despite the limitations imposed by the slow BOLD signal integration in fMRI.
(2) Interpretability: We use voxel-wise and ROI-wise visualization techniques to elucidate the interpretability of our proposed model from a neurobiological perspective.
@@
(3) Comprehensive evaluation: We introduce eight evaluation metrics that comprehensively assess the reconstruction results of our model and all previous models across three dimensions—semantic, structure, and spatiotemporal consistency—on three publicly available video-fMRI datasets. This establishes our work as the first unified benchmark for subsequent researchers. We have released all data and code to facilitate future research: https://github.com/ReedOnePeck/Mind-Animator.
2 Related work
2.1 Reconstructing Videos from Brain Activities
Figure 2: Overview of the video reconstruction paradigms.
The video reconstruction task involves recreating the video frames a subject was viewing based on their brain responses (e.g., fMRI). The challenge in video reconstruction lies in the significant discrepancy between the temporal resolution of fMRI (0.5 Hz) and the frame rate of the stimulus video (30 Hz), making it difficult to model the mapping between fMRI signals and video content.
To tackle this challenge, Nishimoto et al. (2011) reframes video reconstruction as an identification task, using the Motion-Energy model (Adelson & Bergen (1985)) and Bayesian inference to retrieve videos from a predefined library. With the advancement of deep learning, early works by Han et al. (2019), Wen et al. (2018) and Wang et al. (2022), as shown in Figure 2 (a), mapped brain responses to the feature spaces of deep neural networks (DNNs) for end-to-end video reconstruction. To address the scarcity of paired video-fMRI data, Kupershmidt et al. (2022) further advanced this approach by leveraging self-supervised learning to incorporate a large volume of unpaired video data. Although these studies demonstrated the feasibility of reconstructing videos from fMRI signals, the results notably lacked explicit semantic information. As shown in Figure 2 (b), with advancements in multimodal and generative models, Chen et al. (2024), Sun et al. (2024) used contrastive learning to map fMRI signals to the CLIP latent space for semantic decoding, followed by input into a video generation model for reconstruction. This approach produces semantically coherent and smooth videos, but it remains unclear whether the motion information in the reconstructions originates from the fMRI or from the external video data used to train the video generation model.
To address the above issues, we propose Mind-Animator. By independently decoding semantic, structural, and motion information from fMRI signals and inputting them into an inflated image generation model, we ensure that the motion in the reconstructed videos originates solely from the fMRI data.
2.2 Diffusion Models for Video Generation
After significant progress in text-to-image (T2I) generation, diffusion models have drawn interest for text-to-video (T2V) tasks. Ho et al. (2022b) made a breakthrough by introducing 3D diffusion U-Net for video generation, followed by advancements like cascaded sampling frameworks and super-resolution methods (Ho et al. (2022a)), and the integration of temporal attention mechanisms (Singer et al. (2022); Zhou et al. (2022); He et al. (2022b)). However, due to limited paired text-video datasets and high memory demands of 3D U-Nets, alternative approaches have emerged, refining pre-trained T2I models for T2V tasks. Khachatryan et al. (2023) and Wu et al. (2023) introduced techniques like cross-frame attention and inter-frame correlation consideration to adapt T2I models for video generation.
In our work on video reconstruction from fMRI, we avoided pre-trained T2V models to prevent external video data from interfering with motion information decoding from fMRI. As shown in Figure 2 (c), we adapted an inflated T2I model to generate each frame. This ensured that the motion information in the reconstructed videos was derived solely from fMRI decoding, as the generative model had never been exposed to video data.
@@ 3 Methodology
Figure 3: The overall architecture of Mind-Animator, a two-stage video reconstruction model based on fMRI. Three decoders are trained during the fMRI-to-feature stage to disentangle semantic, structural, and motion feature from fMRI, respectively. In the feature-to-video stage, the decoded information is input into an inflated Text-to-Image (T2I) model for video reconstruction.
@@ 3.1 Problem Statement
We aim to decode videos from brain activity recorded with fMRI when healthy participants watch a sequence of natural videos. Let $X$ and $Y$ denote the voxel space and pixel space, respectively. Let $\mathbf{x}_{i}\in\mathbb{R}^{1\times n}$ be the fMRI signal when a video $\mathbf{v}_{i,j}\in\mathbb{R}^{1\times 3\times 512\times 512}$ is presented to the participant, where $n$ is the number of fMRI voxels, $j$ is the frame ID of video $i$ and $i\in[1,N]$, $j\in[1,8]$, with $N$ the total number of videos. Let $Z(k)$ denotes the feature space, $k\in\{semantic,structure,motion\}$. The goal of fMRI-to-feature stage is to train decoders $D(k):X\rightarrow Z(k)$, and the goal of feature-to-video stage is to construct a generation model $G:Z(semantic)\times Z(structure)\times Z(motion)\rightarrow Y$, without introducing motion information from external video data.
@@ 3.2 FMRI-to-feature Stage
Semantic Decoder.
Due to the low signal-to-noise ratio of the fMRI signal $\mathbf{x}_{i}$ and the substantial dimension discrepancy with the text condition $\mathbf{c}_{i}\in\mathbb{R}^{1\times 20\times 768}$ of Stable Diffusion (SD), learning a mapping between them directly is prone to overfitting. Considering both the lower dimensionality ($\mathbf{t}_{i}\ or\ \mathbf{v}_{i}\in\mathbb{R}^{1\times 512}$) and the robust semantic information embedded in the latent space of CLIP (Gao et al. (2020)), and given that CLIP has been shown to outperform various single-modal DNNs in explaining cortical activity (Wang et al. ; Zhou et al. (2024)), we employ bidirectional InfoNCE loss to align the fMRI embedding $\mathbf{f}_{i}$ with the latent space of CLIP (Vit-B/32)$\subseteq\mathbb{R}^{512}$, followed by a two-layer MLP to map it to text condition $\mathbf{c}_{i}$. In this context, $\mathbf{f}_{i}=D_{Semantic}(\mathbf{x}_{i})$, where $D_{Semantic}$ is a three-layer MLP,
$L_{BiInfoNCE}=-\frac{1}{B}\sum_{i=1}^{B}\Big{(}log\frac{exp(s(\mathbf{\hat{z}}_{i},\mathbf{z}_{i})/\tau)}{\sum_{j=1}^{B}exp(s(\mathbf{\hat{z}}_{i},\mathbf{z}_{j})/\tau)}+log\frac{exp(s(\mathbf{\hat{z}}_{i},\mathbf{z}_{i})/\tau)}{\sum_{k=1}^{B}exp(s(\mathbf{\hat{z}}_{k},\mathbf{z}_{i})/\tau)}\Big{)}.$ (1)
where $s(\cdot,\cdot)$ is the cosine similarity, $\mathbf{z}$ and $\mathbf{\hat{z}}$ are the latent representation from two modalities, $B$ is the batch size, and $\tau$ is a learned temperature parameter. Then, given $\mathbf{f}\in\mathbb{R}^{B\times 512}$, $\mathbf{v}\in\mathbb{R}^{B\times 512}$, and $\mathbf{t}\in\mathbb{R}^{B\times 512}$ as the respective embeddings of fMRI, video, and text, the fMRI-vision-language tri-modal loss is:
$L_{Semantic}=\alpha\cdot L_{BiInfoNCE}(\mathbf{f},\mathbf{t})+(1-\alpha)\cdot L_{BiInfoNCE}(\mathbf{f},\mathbf{v}).$ (2)
For stable training, we have also designed the following loss function to bring the fMRI and text embeddings closer together:
$L_{Projection1}=\frac{1}{B}\sum_{i=1}^{B}\|\mathbf{f}_{i}-\mathbf{t}_{i}\|^{2}_{2}.$ (3)
Subsequently, to map the fMRI embedding $\mathbf{f}_{i}$ to the text condition $\mathbf{c}_{i}$ for the purpose of conditioning generative image models, a projection loss is utilized,
$L_{Projection2}=\frac{1}{B}\sum_{i=1}^{B}\|MLP(\mathbf{f}_{i})-\mathbf{c}_{i}\|^{2}_{2}.$ (4)
@@
$L_{Combined}=L_{Projection1}+\lambda_{1}\cdot L_{Semantic}+\lambda_{2}\cdot L_{Projection2}.$ (5)
Structure Decoder.
For a short video clip, it can be assumed that the low-level feature (e.g. size, shape, and color) contained in each frame remains largely consistent with that of the first frame. Consequently, we utilize the token extracted from the first frame by VQ-VAE as structural feature and train the structural decoder (a two-layer MLP) using the standard mean squared error (MSE) loss function. Let $\Phi$ denote the encoder of VQVAE, the structure loss is defined as:
$L_{Structure}=\frac{1}{B}\sum_{i=1}^{B}\|D_{Structure}(\mathbf{f}_{i})-\Phi(\mathbf{v}_{i,1})\|^{2}_{2}.$ (6)
Consistency Motion Generator.
Drawing inspiration from natural language processing, we treat each video frame token as a word embedding and develop an L-layer Transformer-based Consistency Motion Generator (CMG) to implicitly decode dynamic information between consecutive frames.
Figure 4: The architecture of CMG with Temporal Module and fMRI guided Spatial Module.
Handling raw pixels for a video clip $\mathbf{v}_{i}\in\mathbb{R}^{f\times 3\times H\times W}$ in the CMG can be computationally intensive. To overcome this, we follow Stable Diffusion’s approach (Rombach et al. (2022)) by projecting the video into a latent space using a pre-trained VAE tokenizer. After compressing the video clip to $\Phi(\mathbf{v}_{i})\in\mathbb{R}^{f\times 3\times\frac{H}{8}\times\frac{W}{8}}$, it is divided into patches, which are then converted into frame tokens $\Phi_{tok}(\mathbf{v}_{i})\in\mathbb{R}^{P\times d_{token}}$ through an embedding layer.
In the Temporal Module, visible frame tokens $\Phi_{tok}(\mathbf{v}_{i})\in\mathbb{R}^{m\times d_{token}}$ and positional encoding $\mathbf{E_{pos}}\in\mathbb{R}^{m\times d_{token}}$ are jointly input into a Sparse Causal Self-Attention layer to learn inter-frame temporal information. This attention layer incorporates a specially designed Sparse Causal Mask to ensure sparsity between frames. As illustrated in Figure 4, the mask is divided into fixed and random components. The fixed mask ensures that each frame cannot access information from subsequent frames, while the random mask maintains sparsity among visible frames, preventing the model from taking shortcuts (Tong et al. (2022)). During inference, we eliminate the random mask. For other variants of the spatial-temporal module, please refer to Appendix D.2.
In the Spatial Module, to extract spatial information for subsequent frames from fMRI, the embedding of the visible frames $\mathbf{z}_{l}$ serves as the Query, while the fMRI signal $\mathbf{f}$, after passing through an embedding layer, serves as the Key and Value in the cross-attention block, as shown in Eq. (7). Following residual connections and layer normalization, $\mathbf{z}_{l}$ is input into the Feed Forward Network (FFN) to predict the subsequent unseen frame tokens $\hat{\Phi(\mathbf{v}_{i,j})},j\in[m+1,n]$:
$\mathbf{z}_{l}=$ $CrossAttention(\mathbf{Q},\mathbf{K},\mathbf{V}),\quad l=1,2,\ldots,L$ (7)
$\mathbf{Q}=$ $\mathbf{W_{Q}}^{l}\cdot\mathbf{z}_{l},\quad\mathbf{K}=\mathbf{W_{K}}^{l}\cdot Emb(\mathbf{f}),\quad\mathbf{V}=\mathbf{W_{V}}^{l}\cdot Emb(\mathbf{f}),$
$\mathbf{z}_{l}=$ $FFN(LN(\mathbf{z}_{l})+\mathbf{z}_{l-1}).\quad l=1,2,\ldots,L$ (8)
@@
$L_{Consistency}=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=m+1}^{n}\|\hat{\Phi_{tok}(\mathbf{v}_{i,j})}-\Phi_{tok}(\mathbf{v}_{i,j})\|^{2}_{2}.$ (9)